# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 3.1 Observables and derivations
${{\mathfrak{d}}_{l}}^{Y}:=\{{\mathcal{A}}\in{{\mathfrak{d}}_{l}}(Y)\ |\ {\overline{\rm tr}}_{X^{c}}({\mathcal{A}})=0\text{ for any brick }X\subsetneq Y\}.$
Each ${{\mathfrak{d}}_{l}}(Y)$ decomposes as a direct sum ${{\mathfrak{d}}_{l}}(Y)=\bigoplus_{X\subseteq Y}{{\mathfrak{d}}_{l}}^{X}$ over bricks contained in $Y$, and for any brick $X\subseteq Y$ the partial trace ${\overline{\rm tr}}_{Y\backslash X^{c}}$ is the projection onto $\bigoplus_{Z\subseteq X}{{\mathfrak{d}}_{l}}^{Z}\subseteq{{\mathfrak{d}}_{l}}(Y)$. Intuitively, ${{\mathfrak{d}}_{l}}^{Y}$ consists of elements of ${{\mathfrak{d}}_{l}}(Y)$ which are not localized on any brick properly contained in $Y$.
Derivations of ${\mathscr{A}}$ which appear in the physical context are typically only densely defined and have the form
${\mathsf{F}}:{\mathcal{A}}\mapsto\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{Y},{\mathcal{A}}],$
where ${\mathsf{F}}^{Y}\in{{\mathfrak{d}}_{l}}^{Y}$. We are now going to define for every $U\subset{\mathbb{R}}^{n}$ a real Lie algebra ${\mathfrak{D}}_{al}(U)$ which consists of derivations approximately localized on $U$ and such that all ${\mathfrak{D}}_{al}(U)$ have a common dense domain. Moreover, they form a pre-cosheaf of Lie algebras over a certain category of subsets of ${\mathbb{R}}^{n}$.
Definition 3.2.
For any element ${\mathsf{F}}=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$ of $\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ and every $U\subset{\mathbb{R}}^{n}$ we let
@@
and define ${\mathfrak{D}}_{al}(U)\subset\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ as the set of elements ${\mathsf{F}}$ with $\|{\mathsf{F}}\|_{U,k}<\infty$ for all $k\geq 0$.
If $U$ is empty, then for $k>0$ $\|{\mathsf{F}}\|_{U,k}<\infty$ if and only if ${\mathsf{F}}^{Y}=0$ for all $Y\in{\mathbb{B}}_{n}$. Thus ${\mathfrak{D}}_{al}(\emptyset)=0$. We also denote ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})={\mathfrak{D}}_{al}$.
@@
The Hausdorff property follows from the fact that if $\|{\mathsf{F}}\|_{U,k}=0$ for any $k\geq 0$ then ${\mathsf{F}}=0$. To show completeness, suppose $\{{\mathsf{F}}_{m}\}_{m\in\mathbb{N}}\subset{\mathfrak{D}}_{al}(U)$ is Cauchy, i.e. that for any $k\geq 0$ and any $\epsilon>0$ there is an $N\in\mathbb{N}$ such that $m,m^{\prime}\geq N\implies\|{\mathsf{F}}_{m}-{\mathsf{F}}_{m^{\prime}}\|_{U,k}<\epsilon$. For any fixed $Y\in{\mathbb{B}}_{n}$ this implies that $\{{\mathsf{F}}_{n}^{Y}\}$ is Cauchy in ${{\mathfrak{d}}_{l}}^{Y}$ (with the operator norm) and thus converges to a limit ${\mathsf{F}}^{Y}$. Let ${\mathsf{F}}:=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$.
@@
$\leq\epsilon(1+\operatorname{diam}(Y)+d(U,Y))^{-k}+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|.$
@@
Recall that for a nonempty $U\subset{\mathbb{R}}^{n}$ we write $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$. The norms $\|\cdot\|_{U,k}$ obey the following dominance relation.
Lemma 3.2.
Let $U,V$ be subsets of the lattice and suppose that $U\subseteq V^{r}$. Then for any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ we have
@@
Let $Y$ be any subset of the lattice. Then (7) follows from
@@
The above Lemma shows that the space ${\mathfrak{D}}_{al}(U)$ only depends on the asymptotic geometry of the region $U$ in the following sense: if $U\subseteq V^{r}$ and $V\subseteq U^{r}$ for some $r\geq 0$ then ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(V)$ (as subsets of $\prod_{Y}{{\mathfrak{d}}_{l}}^{Y})$ and are isomorphic as Fréchet spaces. In particular, for any non-empty bounded $U\subset{\mathbb{R}}^{n}$, the space ${\mathfrak{D}}_{al}(U)$ coincides with ${\mathfrak{D}}_{al}(\{0\})$.
To relate the spaces ${\mathfrak{D}}_{al}(U)$ to the traditional $C^{*}$-algebraic picture we prove the following:
Proposition 3.2.
Suppose $U\subset{\mathbb{R}}^{n}$ is non-empty and bounded and let ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$. Then the sum $\sum_{X\in{\mathbb{B}}_{n}}{\mathsf{F}}^{X}$ is absolutely convergent and defines a continuous dense embedding of ${\mathfrak{D}}_{al}(U)$ into the subspace of traceless anti-hermitian elements of the algebra ${\mathscr{A}}$.
Proof.
We can assume without loss of generality that $U=\{0\}$. We have
$\sum_{Y\in{\mathbb{B}}_{n}}\|{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}\|_{\{0\},2n+2}\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,0))^{-2n-2},$
and by Lemma 3.1 the above sum is finite. This shows that the map ${\mathfrak{D}}_{al}(\{0\})\to{\mathscr{A}}$ is well-defined and continuous. Its image is dense in the space of traceless anti-hermitian elements of ${\mathscr{A}}$ because it contains the dense subspace ${{\mathfrak{d}}_{l}}$. Finally, to show that it is injective, writing ${\mathcal{A}}:=\sum_{Y\in{\mathbb{B}}_{n}}{\mathsf{F}}^{Y}$ we have
${\mathsf{F}}^{Y}={\overline{\rm tr}}_{Y^{c}}({\mathcal{A}})-\sum_{X\subsetneq Y}{\overline{\rm tr}}_{X^{c}}({\mathcal{A}})$
and so ${\mathcal{A}}=0\implies{\mathsf{F}}^{Y}=0$ for all $Y\in{\mathbb{B}}_{n}$. ∎
Definition 3.3.
We let ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}$ be the image of ${\mathfrak{D}}_{al}(\{0\})$ under the embedding of Prop. 3.2, with the Fréchet topology of ${\mathfrak{D}}_{al}(\{0\})$.
In [11], the algebra ${\mathscr{A}}_{a\ell}$ of almost-local operators was defined as a subspace of ${\mathscr{A}}$ where a countable family of norms similar to (6) takes finite values. Here we equivalently define ${\mathscr{A}}_{a\ell}$ as the set of elements of ${\mathscr{A}}$ whose traceless hermitian and anti-hermitian parts live in ${{\mathfrak{d}}_{al}}$, topologized as ${\mathscr{A}}_{a\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{al}}\otimes{\mathbb{C}})$.
Proposition 3.3.
${\mathscr{A}}_{a\ell}$ is a dense sub-algebra of ${\mathscr{A}}$.
Proof.
${\mathscr{A}}_{a\ell}$ contains all local observables and these are dense in ${\mathscr{A}}$. The fact that ${\mathscr{A}}_{a\ell}$ is closed under multiplication is proven in [11]. ∎
In view of Prop. 3.2 it is natural to make the following definition.
Definition 3.4.
An element ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ is inner iff it is contained in ${\mathfrak{D}}_{al}(U)$ for some bounded $U$.
For any two bounded sets $U,V$, a local observable is strictly localized on both $U$ and $V$ iff it is localized on their intersection, ie. ${\mathscr{A}}(U)\cap{\mathscr{A}}(V)={\mathscr{A}}(U\cap V)$. An analogous relation for the spaces ${\mathfrak{D}}_{al}(U)$ does not hold in general [^7] Indeed, for any two bounded $U,V$ the spaces ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}(V)$ coincide and are nontrivial but if $U$ and $V$ are disjoint then ${\mathfrak{D}}_{al}(U\cap V)=0$., but it does hold if we assume that $U$ and $V$ satisfy the following transversality condition:
Definition 3.5.
@@
Proposition 3.4.
@@
The first inequality is true even without assuming transversality – it follows from Lemma 3.2. For the second, let $Z\in{\mathbb{B}}_{n}$ and choose $x^{*},y^{*}\in Z$ so that $d(x^{*},U)=d(Z,U)$ and $d(y^{*},V)=d(Z,V)$. Then we have
@@
$1+\operatorname{diam}(Z)+d(Z,U\cap V)\leq$
$\hskip 28.45274pt\leq(C+1)(1+\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V))),$ (9)
@@
Proposition 3.5.
For any $U,V\subset{\mathbb{R}}^{n}$ consider the sequence of vector spaces
[…]
@@
$\chi(Y):=\left\{\begin{array}[]{cc}1&\text{ if }d(Y,U)<d(Y,V)\\ 1/2&\text{ if }d(Y,U)=d(Y,V)\\ 0&\text{ if }d(Y,U)>d(Y,V)\end{array}\right..$
@@
$ii)$. Follows immediately from Proposition 3.4. ∎
Next we will show how to endow ${\mathfrak{D}}_{al}(U)$ with the Lie algebra structure.
Proposition 3.6.
Let $U,V\subset{\mathbb{R}}^{n}$. For any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ and ${\mathsf{G}}\in{\mathfrak{D}}_{al}(V)$ the sum
$[{\mathsf{F}},{\mathsf{G}}]^{Z}:=\sum_{X,Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}$ (11)
is absolutely convergent for every $Z\in{\mathbb{B}}_{n}$. The resulting bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] satisfies the Jacobi identity and
@@
for some constant $C>0$ that depends only on $n$.
To prove Proposition 3.6 we will need several lemmas. For any $X,Y\in{\mathbb{B}}_{n}$, define [^8] This is well-defined since the intersection of an arbitrary number of bricks is either empty or a brick, so $X\vee Y$ is the intersection of all bricks containing $X$ and $Y$. the join $X\vee Y\in{\mathbb{B}}_{n}$ as the smallest brick that contains $X$ and $Y$.
Lemma 3.3.
@@
Let $\pi_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ be the projection onto the $i$th coordinate. The following identites hold for any bricks $X,Y\in{\mathbb{B}}_{n}$:
$\pi_{i}(X\vee Y)$ $=\pi_{i}(X)\vee\pi_{i}(Y),$
$\operatorname{diam}(X)$ $=\max_{i=1,\ldots n}\operatorname{diam}(\pi_{i}(X)),$
$d(X,Y)$ $=\max_{i=1,\ldots n}d(\pi_{i}(X),\pi_{i}(Y)).$
@@
Lemma 3.4.
Let $X,Y,Z\in{\mathbb{B}}_{n}$ and let ${\mathsf{F}}\in{{\mathfrak{d}}_{l}}(X)$ and ${\mathsf{G}}\in{{\mathfrak{d}}_{l}}(Y)$. Then $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}=0$ unless $X\cap Y\neq\varnothing$ and $Z\subseteq X\vee Y$.

### v3

@@ 2.1 Observables and derivations
${{\mathfrak{d}}_{l}}^{Y}:=\{{\mathcal{A}}\in{{\mathfrak{d}}_{l}}(Y)\ |\ {\overline{\rm tr}}_{X^{c}}({\mathcal{A}})=0\text{ for any brick }X\subsetneq Y\}.$ (14)
Since each ${{\mathfrak{d}}_{l}}(Y)$ decomposes as a direct sum ${{\mathfrak{d}}_{l}}(Y)=\bigoplus_{X\subseteq Y}{{\mathfrak{d}}_{l}}^{X}$ over bricks contained in $Y$, and since ${{\mathfrak{d}}_{l}}=\bigcup_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}(Y)$, we have a decomposition as the algebraic direct sum ${{\mathfrak{d}}_{l}}=\bigoplus_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$.
Definition 9.
For any $Y\in{\mathbb{B}}_{n}$ define $\pi^{Y}:{{\mathfrak{d}}_{l}}\to{{\mathfrak{d}}_{l}}^{Y}$ as the orthogonal projection.
The following lemma is often useful.
Lemma 10.
There is a function $\mu:{\mathbb{B}}_{n}\times{\mathbb{B}}_{n}\rightarrow{\mathbb{Z}}$ such that for any $X\in{\mathbb{B}}_{n}$ and any ${\mathcal{A}}\in{{\mathfrak{d}}_{l}}$ we have
${\overline{\rm tr}}_{X^{c}}({\mathcal{A}})$ $=\sum_{\begin{subarray}{c}Y\subseteq X\\ Y\in{\mathbb{B}}_{n}\end{subarray}}\pi^{Y}({\mathcal{A}}),$ (15)
$\pi^{X}({\mathcal{A}})$ $=\sum_{\begin{subarray}{c}Y\subseteq X\\ Y\in{\mathbb{B}}_{n}\end{subarray}}\mu(X,Y){\overline{\rm tr}}_{Y^{c}}({\mathcal{A}}).$ (16)
Proof.
The first statement is obvious. The second statement follows from the first one by letting $\mu$ to be the Möbius function [34] of the locally-finite poset $({\mathbb{B}}_{n},\subseteq)$. ∎
Using this lemma and an explicit form of the Möbius function of $({\mathbb{B}}_{n},\subseteq)$ one can show (see Appendix C of [19]):
Lemma 11.
For any $\mathcal{A}\in{{\mathfrak{d}}_{l}}$ and any $Y\in{\mathbb{B}}_{n}$ we have $\|\pi^{Y}(\mathcal{A})\|\leq 4^{n}\,\|\mathcal{A}\|$.
Let $\mathsf{F}$} be a derivation of $\mathscr{A}$} with domain $\mathscr{A}$$_{\ell}$. For any $X\in{\mathbb{B}}_{n}$, the map ${\overline{\rm tr}}_{X^{c}}\circ{\mathsf{F}}|_{{\mathscr{A}}(X)}:{\mathscr{A}}(X)\to{\mathscr{A}}(X)$ is a derivation (this follows from the defining property (5) of the partial trace). We therefore have ${\overline{\rm tr}}_{X^{c}}\circ{\mathsf{F}}|_{{\mathscr{A}}(X)}={\rm ad}_{\mathcal{A}_{X}}$ for some $\mathcal{A}_{X}\in{{\mathfrak{d}}_{l}}(X)$, and we define
${\mathsf{F}}^{X}:=\pi^{X}(\mathcal{A}_{X}).$ (17)
Definition 12.
For any unbounded derivation with domain $\mathscr{A}$$_{\ell}$, the set of local observables $\{{\mathsf{F}}^{X}\}_{X\in{\mathbb{B}}_{n}}$ is the brick decomposition of $\mathsf{F}$}.
The utility of the brick decomposition comes from its uniqueness:
Lemma 13.
Two unbounded derivations of $\mathscr{A}$} with domain $\mathscr{A}$$_{\ell}$ and identical brick decompositions must coincide.
Proof.
An elementary calculation shows that whenever $X\subseteq Y$ we have ${\overline{\rm tr}}_{X^{c}}(\mathcal{A}_{Y})=\mathcal{A}_{X}$, and so we have
$\mathcal{A}_{X}=\bigoplus_{Z\subseteq X}{\mathsf{F}}^{Z}$ (18)
for every $X\in{\mathbb{B}}_{n}$. Thus, for any $X\subseteq Y\in{\mathbb{B}}_{n}$ and any $\mathcal{B}\in{\mathscr{A}}(X)$, the quantity ${\overline{\rm tr}}_{Y^{c}}({\mathsf{F}}(\mathcal{B}))=[\mathcal{A}_{Y},\mathcal{B}]$ is determined by the brick decomposition of $\mathsf{F}$}. Since ${\mathsf{F}}(\mathcal{B})=\lim_{Y\uparrow{\mathbb{R}}^{n}}{\overline{\rm tr}}_{Y^{c}}({\mathsf{F}}(\mathcal{B}))$, it follows that ${\mathsf{F}}(\mathcal{B})$ is determined by the brick decomposition of $\mathsf{F}$}. ∎
Lemma 13 allows us to identify an unbounded derivation $\mathsf{F}$} with its set of brick components $\{{\mathsf{F}}^{X}\}_{X\in{\mathbb{B}}_{n}}$. In what follows, we will refer to a derivation and its brick decomposition interchangeably  [^3] However, note that not every element of $\prod_{X\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{X}$ defines a derivation that is convergent on $\mathscr{A}$$_{\ell}$..
Example 14.
If ${\mathsf{F}}={\mathsf{F}}_{\Phi}$ for some interaction $\Phi$ as in definition 4, then we have
${\mathsf{F}}^{X}=\sum_{Y\in\mathcal{P}(\Lambda)}\pi^{X}(\Phi(Y)),\quad\forall X\in{\mathbb{B}}_{n}.$ (19)
The sum is absolutely convergent by (6) and Lemma 11.
Uniformly almost-local derivations
As Examples 5 and 6 show, physically relevant derivations are unbounded and cannot be continuously extended to the whole quasilocal algebra $\mathscr{A}$}. To apply algebraic tools, one would like to define a space of derivations which is large enough to cover examples of interest and is a topological Lie algebra.
We will use the brick decomposition to define such a class of derivations:
Definition 15.
For any element ${\mathsf{F}}=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$ of $\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ we let
$\|{\mathsf{F}}\|_{k}:=\sup_{Y\in{\mathbb{B}}_{n}}\|{\mathsf{F}}^{Y}\|(1+\operatorname{diam}(Y))^{k}.$ (20)
Proposition 16.
Let ${\mathsf{F}}\in\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ and $\|{\mathsf{F}}\|_{k}<\infty$ for all $k\geq 0$. Then there exists an interaction $\Phi$ such that the brick components of ${\mathsf{F}}_{\Phi}$ are given by $\mathsf{F}$}.
Proof.
For any $X\in{\mathcal{P}}(\Lambda)$ let
$\Phi(X)=\left\{\begin{array}[]{cc}{\mathsf{F}}^{\lceil X\rceil}&\text{ if $\lceil X\rceil\cap\Lambda=X$}\\ 0&\text{ otherwise}\end{array}\right$. (23)
where $\lceil X\rceil\in{\mathbb{B}}_{n}$ is the meet of all bricks containing $X$. Clearly, $\Phi(X)\in{{\mathfrak{d}}_{l}}(X)$. Also, for any $x\in\Lambda$ we have
$\sum_{\begin{subarray}{c}X\in{\mathcal{P}}(\Lambda)\\ X\ni x\end{subarray}}\|\Phi(X)\|\leq\sum_{\begin{subarray}{c}Y\in{\mathbb{B}}_{n}\\ Y\ni x\end{subarray}}\|{\mathsf{F}}^{Y}\|\leq\|{\mathsf{F}}\|_{2n+2}\sum_{\begin{subarray}{c}Y\in{\mathbb{B}}_{n}\\ Y\ni x\end{subarray}}(1+\operatorname{diam}(Y))^{-2n-2}\leq\infty,$ (24)
where in the last step we used Lemma 8. Thus, $\Phi$ is a well-defined interaction. Finally, using (19) it is easy to check that $\left({\mathsf{F}}_{\Phi}\right)^{Y}={\mathsf{F}}^{Y}$ for any $Y\in{\mathbb{B}}_{n}$. ∎
Definition 17.
Define the space ${\mathfrak{D}}_{al}$ as the set of elements ${\mathsf{F}}\in\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ with $\|{\mathsf{F}}\|_{k}<\infty$ for all $k\geq 0$.
In [19] ${\mathfrak{D}}_{al}$ was called the space of uniformly almost local derivations (UAL derivations). From the physical viewpoint, UAL Hamiltonians correspond to interactions between sites of $\Lambda$ which decay faster than any power of distance.
Example 18.
Let $\Lambda={\mathbb{Z}}^{n}\subseteq{\mathbb{R}}^{n}$. The derivation (8) corresponds to ${\mathsf{Q}}^{Y}$ being nonzero only for bricks of zero diameter, in which case ${\mathsf{Q}}^{\{x\}}={\mathsf{q}}_{x}$. If ${\mathsf{q}}_{x}$ is uniformly bounded, such $\mathsf{Q}$} is a UAL derivation. If $\Lambda$ is an arbitrary lattice in ${\mathbb{R}}^{n}$, a derivation of the form (8) has ${\mathsf{Q}}^{Y}$ nonzero only for $\operatorname{diam}Y\leq 1$. With our assumptions on $\Lambda$, $\mathsf{Q}$} is a UAL derivation whenever ${\mathsf{q}}_{x}$ is uniformly bounded.
Example 19.
Consider a system of fermionic particles as in Example 6 with all $H_{x}$ isomorphic to a fixed finite-dimensional Hilbert space $H$. If the particles are non-interacting, the corresponding Hamiltonian is described by a derivation of the form
${\mathsf{F}}(h):{\mathcal{A}}\mapsto\sum_{x,y\in\Lambda}\left[\sum_{\alpha,\beta}c^{*}_{\alpha,x}c_{\beta,y}h_{\alpha\beta}(x,y),{\mathcal{A}}\right],$ (25)
where $h_{\alpha\beta}(x,y)$ are complex numbers. One may regard these numbers as elements of an infinite square matrix whose rows and columns are labeled by a pair $(\alpha,x)$, or as matrix elements of an operator on $\ell^{2}(\Lambda,H)$. $\mathsf{F}$} is anti-hermitian iff this matrix is anti-hermitian. Equivalently, one can view the numbers $h_{\alpha\beta}(x,y)$ as defining a function $h:\Lambda\times\Lambda\rightarrow\mathrm{Hom}_{\mathbb{C}}(H,H)$. An anti-hermitian ${\mathsf{F}}(h)$ is a UAL derivation if $h$ is bounded and $h(x,y)=O(|x-y|^{-\ell})$ for all $\ell\geq 0$. The charge derivation $\mathsf{Q}$} from Example 6 also has the form (25), with $h(x,y)=\delta_{xy}\in\mathrm{Hom}_{\mathbb{C}}(H,H)$ for all $x,y\in\Lambda$.
The functions $\|\cdot\|_{k}$ are norms. We show in the following that ${\mathfrak{D}}_{al}$ equipped with these norms is a Fréchet -Lie algebra. But first we generalize the definition of ${\mathfrak{D}}_{al}$ by defining for every $U\subseteq{\mathbb{R}}^{n}$ a real Lie algebra ${\mathfrak{D}}_{al}(U)$ which consists of derivations approximately localized on $U$.
Definition 20.
For any element ${\mathsf{F}}=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$ of $\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ and every $U\subseteq{\mathbb{R}}^{n}$ we let
@@
Definition 21.
Define ${\mathfrak{D}}_{al}(U)\subseteq\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ as the set of elements $\mathsf{F}$} with $\|{\mathsf{F}}\|_{U,k}<\infty$ for all $k\geq 0$.
If $U$ is empty, then for $k>0$ $\|{\mathsf{F}}\|_{U,k}<\infty$ if and only if ${\mathsf{F}}^{Y}=0$ for all $Y\in{\mathbb{B}}_{n}$. Thus ${\mathfrak{D}}_{al}(\emptyset)=\{0\}$. On the other hand we have ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})={\mathfrak{D}}_{al}$.
@@
The Hausdorff property follows from the fact that if $\|{\mathsf{F}}\|_{U,k}=0$ for any $k\geq 0$ then ${\mathsf{F}}=0$. To show completeness, suppose $\{{\mathsf{F}}_{m}\}_{m\in\mathbb{N}}\subseteq{\mathfrak{D}}_{al}(U)$ is Cauchy, i.e. that for any $k\geq 0$ and any $\epsilon>0$ there is an $N\in\mathbb{N}$ such that $m,m^{\prime}\geq N\implies\|{\mathsf{F}}_{m}-{\mathsf{F}}_{m^{\prime}}\|_{U,k}<\epsilon$. For any fixed $Y\in{\mathbb{B}}_{n}$ this implies that $\{{\mathsf{F}}_{m}^{Y}\}$ is Cauchy in ${{\mathfrak{d}}_{l}}^{Y}$ (with the operator norm) and thus converges to a limit ${\mathsf{F}}^{Y}$. Let ${\mathsf{F}}:=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$.
@@
$\leq\epsilon(1+\operatorname{diam}(Y)+d(U,Y))^{-k}+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|.$ (27)
@@
Definition 23.
Let $U$ be a non-empty subset of ${\mathbb{R}}^{n}$. For any $r\geq 0$ the $r$-thickening of $U$, denoted $U^{r}$, is $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$.
The norms $\|\cdot\|_{U,k}$ obey the following dominance relation.
Lemma 24.
Let $U,V$ be subsets of ${\mathbb{R}}^{n}$ and suppose that $U\subseteq V^{r}$. Then for any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ we have
@@
Let $Y\in{\mathbb{B}}_{n}$. Then (28) follows from
@@
The above Lemma shows that the space ${\mathfrak{D}}_{al}(U)$ only depends on the asymptotic geometry of the region $U$. More precisely, we have
Corollary 25.
If $U\subseteq V^{r}$ and $V\subseteq U^{r}$ for some $r\geq 0$ then ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(V)$ (as subsets of $\prod_{Y}{{\mathfrak{d}}_{l}}^{Y})$ and are isomorphic as Fréchet spaces. In particular, for any non-empty bounded $U\subseteq{\mathbb{R}}^{n}$, the space ${\mathfrak{D}}_{al}(U)$ coincides with ${\mathfrak{D}}_{al}(\{0\})$.
Behaviour under intersection
For any two bounded sets $U,V$, a local observable is strictly localized on both $U$ and $V$ iff it is localized on their intersection, i.e. ${\mathscr{A}}(U)\cap{\mathscr{A}}(V)={\mathscr{A}}(U\cap V)$. An analogous relation for the spaces ${\mathfrak{D}}_{al}(U)$ does not hold in general [^4] Indeed, for any two bounded $U,V$ the spaces ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}(V)$ coincide and are nontrivial but if $U$ and $V$ are disjoint then ${\mathfrak{D}}_{al}(U\cap V)=0$., but it does hold if we assume that $U$ and $V$ satisfy the following transversality condition:
Definition 26.
@@
Example 27.
  • (i)
    For any $\subseteq{\mathbb{R}}^{n}, UU$ and ${\mathbb{R}}^{n}$ are transverse.
  • (ii)
    If $U\cap V=\emptyset$, then $U$ and $V$ are not transverse.
  • (iii)
    If $U$ and $V$ are two intersecting straight lines in ${\mathbb{R}}^{2}$, then $U$ and $V$ are transverse.
  • (iv)
    If $U=\{(x,0)\in{\mathbb{R}}^{2}\}$ and $V=\{(x,x^{1/3})\in{\mathbb{R}}^{2}\}$, then $U$ and $V$ are not transverse.
Transversality of two regions $U,V$ implies the following important compatibility between their Fréchet norms:
Proposition 28.
@@
The first inequality is true even without assuming transversality – it follows from Lemma 24. For the second, note that we can assume without loss of generality that $U$ and $V$ are closed. Let $Z\in{\mathbb{B}}_{n}$ and choose $x^{*},y^{*}\in Z$ so that $d(x^{*},U)=d(Z,U)$ and $d(y^{*},V)=d(Z,V)$. Then we have
@@
$1+\operatorname{diam}(Z)+d(Z,U\cap V)\leq(C+1)(1+\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V))),$ (33)
@@
Proposition 29.
For any $U,V\subseteq{\mathbb{R}}^{n}$ consider the sequence of vector spaces
${\mathfrak{D}}_{al}(U\cap V)\xrightarrow{\alpha}{\mathfrak{D}}_{al}(U)\oplus{\mathfrak{D}}_{al}(V)\xrightarrow{\beta}{\mathfrak{D}}_{al}(U\cup V)\to 0,$ (34)
@@
$\chi(Y):=$ $\left\{\begin{array}[]{cc}1&\text{ if }d(Y,U)<d(Y,V)\\ 1/2&\text{ if }d(Y,U)=d(Y,V)\\ 0&\text{ if }d(Y,U)>d(Y,V)\end{array}\right$. (38)
@@
ii)ii). Follows immediately from Proposition 28. ∎
Next we will show how to endow ${\mathfrak{D}}_{al}(U)$ with the structure of a Lie algebra.
Proposition 30.
Let $U,V\subseteq{\mathbb{R}}^{n}$. For any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ and ${\mathsf{G}}\in{\mathfrak{D}}_{al}(V)$ the sum
${\mathsf{H}}^{Z}:=\sum_{X,Y\in{\mathbb{B}}_{n}}\pi^{Z}\left([{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]\right)$ (39)
is absolutely convergent for every $Z\in{\mathbb{B}}_{n}$. Here $\pi^{Z}$ is the orthogonal projection onto ${{\mathfrak{d}}_{l}}^{Z}$ in ${{\mathfrak{d}}_{l}}=\oplus_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$. The brick components ${\mathsf{H}}^{Z}$ define an element of $\prod_{Z\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Z}$ which we call ${\mathsf{F}},{\mathsf{G}}]$, which satisfies
@@
for some constant $C>0$ that depends only on $n$. The resulting bracket $\cdot,\cdot]:{\mathfrak{D}}_{al}(U)\times{\mathfrak{D}}_{al}(V)\to{\mathfrak{D}}_{al}(U)$ is bilinear, skew-symmetric, and satisfies the Jacobi identity.
To prove Proposition 30 we will need several lemmas.
Lemma 31.
@@
Let $\pi_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ be the projection onto the $i$th coordinate. The following identities hold for any bricks $X,Y\in{\mathbb{B}}_{n}$:
$\pi_{i}(X\vee Y)$ $=\pi_{i}(X)\vee\pi_{i}(Y),$ (43)
$\operatorname{diam}(X)$ $=\max_{i=1,\ldots n}\operatorname{diam}(\pi_{i}(X)),$ (44)
$d(X,Y)$ $=\max_{i=1,\ldots n}d(\pi_{i}(X),\pi_{i}(Y)).$ (45)
@@
Lemma 32.
Let $X,Y,Z\in{\mathbb{B}}_{n}$ and let ${\mathcal{A}}\in{{\mathfrak{d}}_{l}}(X)$ and ${\mathcal{B}}\in{{\mathfrak{d}}_{l}}(Y)$. Let ${\mathcal{A}}^{X}$ denote the component of $\mathcal{A}$} in ${{\mathfrak{d}}_{l}}^{X}$ and ${\mathcal{B}}^{Y}$ denote the component of $\mathcal{B}$} in ${{\mathfrak{d}}_{l}}^{Y}$. Then the ${{\mathfrak{d}}_{l}}^{Z}$-component of ${\mathcal{A}}^{X},{\mathcal{B}}^{Y}]$ vanishes unless $X\cap Y\neq\varnothing$ and $Z\subseteq X\vee Y$.